Assume that y vaires inversly as x.If y=8 when x=2.5,find x when y=2.
Answer by venugopalramana(3286)

SEE THE FOLLOWING .YOU CAN DO YOUR PROBLEM THEN..YOU WILL GET Y=8*2.5/X=20/X
y waries inversely as x...THIS MEANS Y=K/X WHERE K IS A CONSTANT
and y=18 when x=3.SUBSTITUTING IN ABOVE EQN.,WE GET ..18=K/3..OR..K=18*3=54
Find the constant of variation....=54
Then write an equation for the relation....Y=54/X